[5 Marks]

e) What would be the Price Elasticity of Demand for tickets from Lusaka to Livingstone if the price rises again from 250 kwacha to 300 kwacha? [5 Marks]

f) Explain why the answer in part d) and part e) are different.


1
Expert's answer
2021-09-12T16:30:31-0400

e.

The demand equation given in the question is Qd ​=2000−5P

We can calculate the change in quantity demanded using the equation

When P = 250 "Q_{d}" = 2000- 5(250)

"Q_{d}" = 750 ("Q_{1}" )

When P = 300 "Q_{d}" = 2000- 5(300)

"Q_{d}" = 500 ("Q_{2}\n\u200b" )

Price elasticity of demand "=\frac{\frac{(Q_2-Q_1)}{(Q_1+Q_2)/2}}{\frac{(P_2-P_1)}{(P_2+P_1)/2}}"


Where

"Q_{1}" = 750

"Q_{2}" = 500

"P_{1}" = 250

"P_{2}" = 300


PED = "=\frac{\frac{(500- 750)}{(750+500)/2}}{\frac{(300-250)}{(300+250)/2}}"


PED = -2.2


f.

From question part d the PED was -1.3 therefore PED was elastic as the absolute value of elasticity was more than 1. (1.3>1).

In this part e the elasticity is -2.2. The absolute value is 2.2 which greaters than 1. The difference is that the elasticity of demand in part e is more than the one in part d in absolute terms.

The difference has been necessitated by a huge increase in part e which is way above the increase that was experienced in part d.